OpenStudy (anonymous):

Simplify. Square root -25 square root -9 plz help!

OpenStudy (anonymous):

square root of -25 equals \[ \pm 5i\] and square root of -9 equals \[\pm 3i\]

OpenStudy (anonymous):

So is the answer 3i?

OpenStudy (anonymous):

\[\sqrt{-25}= \sqrt{25(-1)}= \sqrt{25i^2}= \sqrt{(5i)^2}=5i\\\\ \sqrt{-9}= \sqrt{9(-1)}= \sqrt{9i^2}= \sqrt{(3i)^2}=3i\\\\ 5i*3i=15i^2=-15 \]

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

kk